OpenStudy (anonymous):

28=8x+12-7x If I have to add 8x and 7x, where would the number be at?

zepdrix (zepdrix):

You mean the +12 term?

zepdrix (zepdrix):

Keep in mind, you are subtracting 7x's from 8x's. Not adding them :D

OpenStudy (anonymous):

is the problem even possible?? sorry, i cant think tight now D:

OpenStudy (anonymous):

right*

OpenStudy (anonymous):

im just so confused with this

zepdrix (zepdrix):

Keep in mind with addition, you can add/subtract in any order you want. So: 28=8x+12-7x Is the same thing as: 28=8x-7x+12 So let's deal with the x's first. If you have 8 of something, and you take 7 of those things away, how many things are left? XD

OpenStudy (anonymous):

1 lol

zepdrix (zepdrix):

Good, and our "thing" is x :) so let's write it that way, we have 1 x left over 28=8x-7x+12 28=1x+12 28=x+12 Look ok so far? :) now we want to get x alone, hmmmm

OpenStudy (anonymous):

-12 from both sides?

zepdrix (zepdrix):

Yes good good :)